OpenStudy (anonymous):

Real Analysis... Prove directly from the definition that lim n^2/n+1 =1

OpenStudy (anonymous):

limit as n goes to what?

OpenStudy (anonymous):

If there's no error in that question, it would be a really weird limit...

OpenStudy (anonymous):

Is it \[ \frac{n^2}{n} + 1 \] or \[\frac{n^2}{n+1} \] ?

OpenStudy (anonymous):

i bet it is \[\lim_{n\to \infty}\frac{n^2}{n^2+1}\]

OpenStudy (anonymous):

\[\lim \frac{ n^{2} }{ n^{2}+1 }=1\] This is similar to something we did in class so I modeled it after that but I don't know if its right. This is what I have: PROOF: Let eplison >0. \[\left| \frac{ n^{2} }{ n^{2}+1 }- 1 \right|= \left| \frac{ -n-1 }{ n^{2}+1 } \right|= \frac{ n+1 }{ n^{2}+1 }\] Replace n+1 with something larger; 2n \[\frac{ n+1 }{ n^{2}+1 }<\frac{ 2n }{ n^{2}+1 }\] then replace n^2+1 by something smaller; n^2 \[\frac{ 2n }{ n^2 +1} <\frac{ 2n }{ n^2 }=\frac{ 2 }{ n }\] Thus, \[\left| \frac{ n^2 }{ n^2 +1}-1 \right|<\frac{ 2 }{ n }\] Now, choose N, so N>2/epsilon. Then, if \[n \ge N\], \[\left| \frac{ n^2 }{ n^2 +1 }-1 \right| <\frac{ 2 }{ n }<\frac{ 2 }{ N }< \epsilon \]

OpenStudy (anonymous):

I'm horrible with proofs and I'm really not confident that this is right....

OpenStudy (anonymous):

what is the original question?

OpenStudy (anonymous):

\[\frac{n^2}{n^2+1}-1=\frac{n^2-(n^2+1)}{n^2+1}=\frac{-1}{n^2+1}\]

OpenStudy (anonymous):

2a

OpenStudy (dape):

Well, for one \[ \left|\frac{n^2}{n^2+1}-1\right| \neq \frac{n+1}{n^2+1} \] but \[ \left|\frac{n^2}{n^2+1}-1\right| = \frac{1}{n^2+1} \]

OpenStudy (anonymous):

Your proof looks good in principle but the algebra has an error, as the gentleman above has pointed out.

OpenStudy (anonymous):

so let \(\epsilon>0\) and you want \(\frac{1}{n^2+1}<\epsilon\) so take \(n^2+1>\frac{1}{\epsilon}\) which you can do by whatever that thing is called that says the natural numbers are not bounded above

OpenStudy (anonymous):

or if you want to be annoyingly pedantic, take \[n>\sqrt{\frac{1}{\epsilon}-1}\]

OpenStudy (anonymous):

Ok... I think I get this now. Im going to go work on it. Thank you all for your help!!!